The Ka in my textbook for HSCN is .13.Calculate the pH of a solution prepared by (This problem requires values in your textbooks specific appendices, which you ca
Answer #1

0.13) becaus dissociate in Its pH HSCN is a weak acid (ka= 0.13) because it does not Completely dissociate into aqueous solut. The concentration of this HSCN Solution = 0.09037(M). (M= molarity concentration unit) PH of 0.09037(M) HSCN solution ... H- Note - If we change the volume of s, concentration solution to the new volume of V₂ by any means. (like addition of more sopH of this solution pH-t (pka - loge). here e= 0.009037(1). St (0.886 - log 0.009037) .zţ [0.886 - (- 2.044)] 2:10 465. PhpH of this solution, pH=[0.886-log(9.037x16 53] = $(0.886-(-4.044)] = 2.465. 2. pH of the solution Ľ diluting some of solutio